OpenStudy (surana):

Solve the quadratic equation 5x^2 -125 = 0.

OpenStudy (surana):

The ^2 represents Exponent of 2.

OpenStudy (michele_laino):

hint: \[5{x^2} - 125 = 5\left( {{x^2} - {5^2}} \right) = 5\left( {x - 5} \right)\left( {x + 5} \right)\]

OpenStudy (surana):

That'll help, thanks.

OpenStudy (michele_laino):

thanks!

OpenStudy (surana):

You're welcome.

OpenStudy (michele_laino):

thanks! :)

OpenStudy (surana):

I'm just going to solve it on a piece of paper and then come back with the answer.

TheSmartOne (thesmartone):

Not a lot more work can be done on paper. :P You just now have to set each one to 0, to find the roots (or the value of x) :)

OpenStudy (anonymous):

Not a lot more work can be done on paper. :P You just now have to set each one to 0, to find the roots (or the value of x) :)

OpenStudy (surana):

Set to zero?

TheSmartOne (thesmartone):

\(\sf 5(x+5)(x-5)\) And then to find the value of x: \(\sf (x+5)=0\) and \(\sf (x-5)=0\) And solve for x in both equations, and you should have two values for x as your answer :)

OpenStudy (anonymous):

5(x+5)(x−5) And then to find the value of x: (x+5)=0 and (x−5)=0 And solve for x in both equations, and you should have two values for x as your answer :)

OpenStudy (surana):

So I have to multiply x+5 by 5 and then put five on the side of x and then divide?

TheSmartOne (thesmartone):

No, you just subtract 5 on both sides. And for the other equation, you add 5 to both sides. :)

OpenStudy (anonymous):

No, you just subtract 5 on both sides. And for the other equation, you add 5 to both sides. :)

OpenStudy (surana):

Thank you, I will work that out then.

OpenStudy (surana):

I got -5 and +5.

TheSmartOne (thesmartone):

Correct. :) So now if you wanted to test out those answers, then plug in x=5 and x=-5 into the original equation, and it would equal to 0! :)

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

uh yae

OpenStudy (surana):

So one of those two is the answer, I'll check, then.

OpenStudy (michele_laino):

both of those numbers are the answer

TheSmartOne (thesmartone):

^

OpenStudy (surana):

The thing is that the answers I'm allowed is that only one of those can be chosen. I still don't know which one to choose.

OpenStudy (surana):

I think I'll go with -5.

TheSmartOne (thesmartone):

Do you have answer choices? If the answer box is a blank then try putting `5, -5`

OpenStudy (poopsiedoodle):

is there a choice for \(\pm\)5?

OpenStudy (surana):

Yeah, there is.

OpenStudy (michele_laino):

is x a length?

OpenStudy (surana):

Does that mean both positive and negative?

OpenStudy (poopsiedoodle):

yes, it does

OpenStudy (surana):

Then I'll go with that one instead.

TheSmartOne (thesmartone):

:)

OpenStudy (surana):

Thank you all!

OpenStudy (surana):

I'll close the question now.

OpenStudy (poopsiedoodle):

good luck with your future math things :*